OpenStudy (wavykelp):

Which number is between 0.5 and 5/8? Can you please help me understand this?

OpenStudy (anonymous):

O.5 5/8 are the same number You cant have a number inbetween two numbers that are equil

OpenStudy (shamil98):

0.5 = 1/2 5/8 = is a bit bigger than one half to be exact it is 0.625 so between 0.5 and 0.625

OpenStudy (wavykelp):

uhhm 0.5 = 4/8 doesnt it?

OpenStudy (shamil98):

Those aren't the same numbers..

OpenStudy (shamil98):

Yes, @WavyKelp

OpenStudy (anonymous):

Oh sorry my math is off XD

OpenStudy (anonymous):

and im not wearing my glasses

OpenStudy (wavykelp):

well anyway i need some help with this

OpenStudy (anonymous):

So yes hes right